OpenStudy (anonymous):

Solve the differential equation dy/dx=14x^2y^2 with the condition y(2)=4 the solution to the equation is y=

OpenStudy (anonymous):

its a separable differential equation, \[\frac{dy}{dx}=14x^2y^2\]\[\frac{dy}{y^2}=14x^2dx\]

OpenStudy (anonymous):

and what do we do next?

OpenStudy (espex):

Integrate both sides.

OpenStudy (anonymous):

can you show me step by step?

OpenStudy (espex):

How about you show me your steps and I will help you when you stumble.

OpenStudy (anonymous):

im not sure how to start, thats why i asked